# Diff-PCR: Diffusion-Based Correspondence Searching in Doubly Stochastic Matrix Space for Point Cloud Registration

### 03 Jan 2024 v2

@@ I Introduction
  • To the best of our knowledge, we are the first to utilize the diffusion model in the doubly stochastic matrix space for iteratively searching optimal matching matrix solutions using the reverse sampling process. This matching matrix diffusion model can be seamlessly integrated as a module in any other 2D-2D, 2D-3D, and 3D-3D registration problems.
@@ III-A Preliminaries
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}=\prod_{s=1}^{t}\alpha_{s}$. Based on equation Eqn.2 and rewritten form of transition kernel $q(x_{t}|x_{t-1})=q(x_{t}|x_{t-1},x_{0})$ (due to Markov property), we can derive the ELBO:
$\begin{split}&logp(x_{0})\geq-L_{vb}=\\ &\underbrace{\mathbb{E}_{q(x_{1}|x_{0})}[logp_{\theta}(x_{0}|x_{1})]}_{reconstruction\ term}\\ &-\underbrace{D_{KL}(q(x_{T}|x_{0}||p(x_{T})))}_{prior\ matching\ term}\\ &-{\Sigma_{T=2}^{T}}\underbrace{\mathbb{E}_{q(x_{t}|x_{0})}[D_{KL}(q(x_{t-1}|q_{x_{t}},x_{0}))||p_{\theta}(x_{t-1}|x_{t})]}_{denoising\ matching\ term}\end{split}$ (5)
@@
Based on Eqn.3, Eqn.4, and Eqn.6 and utilize the Bayes rule, the denoising step can be derived as:
$\begin{split}&q(x_{t-1}|x_{t},x_{0})\\ &=\frac{q(x_{t}|x_{t-1},x_{0})q(x_{t-1}|x_{0})}{q(x_{t}|x_{0})}\\ &\propto N(x_{t-1};\underbrace{\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})x_{t}+\sqrt{\bar{\alpha}_{t-1}}(1-\alpha_{t})x_{0}}{1-\bar{\alpha}_{t}}}_{\mu_{q}(x_{t},x_{0})},\\ &\underbrace{\frac{(1-\alpha_{t})(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\textbf{I}}_{\Sigma_{q}(t)}).\end{split}$ (7)
By some derivations[43], the optimization of denoising matching term (in Eqn.5) boils down to learning a network $g_{\theta}(x_{t})$ to predict the ground truth $x_{0}$ (i.e., in $\mu_{q}(x_{t},x_{0})$) from an arbitrarily diffused version $x_{t}$:
@@
Figure 2: Overview of our matching matrix diffusion model. $\bigoplus$ mean 3D point coordinates and position encoding are both utilized as input. $\bigotimes$ means only 3D point utilized. The input points $\hat{P},\hat{Q}$ along with their corresponding point features $F^{\hat{P}},F^{\hat{Q}}$ remain fixed throughout the entire reverse sampling process after being output from the KPConv backbone. These inputs undergo transformation through the warping operation and denoising transformer at each denoising step. The forward diffusion process is modeled by the Gaussian transition kernel $q(E^{t}|E^{t-1})$ which has a closed form $q(E^{t}|E^{0})$. The denoising model $g_{\theta}(E^{t})$ learns a reverse denoising gradient that points to the target solution $E^{0}$. During inference in the reverse sampling process, we utilize the predicted $\hat{E}_{0}$ and Eqn.(6,14) to sampling $E^{t-1}$.
III-B Problem Formulation
@@ III-C Framework Overview
Our framework consists of a KPConv[9] feature backbone and one diffusion model[21]. The backbone takes source point clouds $P$ and target $Q$ as input and performs three downsamplings and two upsamplings to obtain the downsampled super points $\hat{P}$ and $\hat{Q}$, along with their associated features associate features $F^{\hat{P}}\in\mathbb{R}^{N{\times}d}$ and $F^{\hat{Q}}\in\mathbb{R}^{M{\times}d}$. The initilized matching matrix $E^{T}$ is computed by the inner product of $F^{\hat{P}}$ and $F^{\hat{Q}}$. We then employ the denoising module $g_{\theta}$ (III-E) to reverse sample the target matching matrix $E^{0}$.
@@ III-D Diffusion Model in Doubly Stochastic Matrix space
and it is defined only depend on $\alpha_{t}$,$\bar{\alpha}_{t}$,and $E^{0}$ (see Eqn.7).
@@
Optimization loss. Inspired by [46] and Eqn.8, we replace the ”denoising matching term” in Eqn.5 by a simplified version:
$L_{simple}=-\mathbb{E}_{q(\hat{E}^{0})}\left[\Sigma_{t-1}^{T}\left(\frac{T-t+1}{T}\right)\mathbb{E}_{q(E^{t}|\hat{E}^{0})}logp_{\theta}(\hat{E}^{0}|E^{t})\right]$ (15)
where the reweighting term $\frac{T-t+1}{T}$ means that the loss at time steps close to T are weighted less than earlier steps.
@@ III-E2 Weighted SVD: $\mathbf{soft\_procrustes(\cdot,\cdot,\cdot)}$
$H=\Sigma_{(i,j)\in K}\tilde{E}(i,j)\hat{p}_{i}\hat{q}_{j}^{T},\ H=U{\Sigma}V^{T}$ (17)
@@
$\mathbf{t}=\frac{1}{|K|}\left({\Sigma_{(i,\cdot)\in K}\ \hat{p}_{i}}-\mathbf{R}{\Sigma_{(\cdot,j)\in K}}\hat{q}_{j}\right)$ (19)
@@ III-E3 Warping Function: $\mathbf{warping(\cdot,\cdot,\cdot)}$
In this paper, we utilize the rigid warping for both rigid and deformable registration cases to fastly demonstrate our idea. With the predicted correspondences and their associated local rigid transformation assumptions, we can perform nearest neighbor interpolation in the predicted correspondences to interpolate the flow of any point in $\hat{P}$. For the deformation warping, we actually can build a deformation graph[48, 49] to conduct deformable warping[50, 13] for the source point cloud $\hat{P}$. We leave the deformable warping integration in future work.
Remarks. When extending this matching matrix denoising function to other registration problems, such as 2D-2D or 2D-3D registration, you may encounter different types of warping functions beyond simple translation and rotation. Affine transformation is a common choice when dealing with more complex deformations, as it allows for scaling, shearing, and non-uniform transformations.
@@ III-E4 Denoising Transformer: $f_{\theta}(\cdot,\cdot,\cdot,\cdot,\cdot,\cdot)$
Attention layer. Following [5], we integrate the rotary position encoding $\Theta(\cdot)$ (see III-E6) in the attention layers. Specifically, in the self-attention, the $q,k,v$ are computed as:
@@
III-E6 Rotary Position Encoding $\Theta(\cdot)$
Since KPConv [9] is a translation invariant backbone, to avoid the asymmetric or globally repetitive ambiguities, following [5], we utilize the rotary positional encoding $\Theta(\cdot)$ to give the point position embedding. Given a 3D point $\hat{p}_{i}\in\mathbb{R}^{3}$ and associate feature $f^{\hat{p}_{i}}\in\mathbb{R}^{d}$, the position encoding is defined as following:
$\tiny\mathbb{PE}(p_{i},f^{\hat{p}_{i}})=\Theta(p_{i})f^{\hat{p}_{i}}=\begin{pmatrix}M_{1}&&&&\\ &M_{2}&&&\\ &&M_{3}&&\\ &&&M_{4}&\\ &&&&M_{d/6}\\ \end{pmatrix}f^{\hat{p}_{i}}$ (29)
$\tiny M_{k}=\begin{pmatrix}cos(x\theta_{k})&-sin(x\theta_{k})&0&0&0&0\\ sin(x\theta_{k})&cos(x\theta_{k})&0&0&0&0\\ 0&0&cos(y\theta_{k})&-sin(y\theta_{k})&0&0\\ 0&0&sin(y\theta_{k})&cos(y\theta_{k})&0&0\\ 0&0&0&0&cos(z\theta_{k})&-sin(z\theta_{k})\\ 0&0&0&0&sin(z\theta_{k})&cos(z\theta_{k})\\ \end{pmatrix}$ (36)
where $\theta_{k}=\frac{1}{10000^{6(k-1)/d}}$ and $k\in[1,2,...,d/6]$.
@@ III-F Framework Training
Our framework utilize a KPConv[9] branch to give the downsampled the superpoints $\hat{P}$ and $\hat{Q}$ and the asccociated features $F^{\hat{P}}$ and $F^{\hat{Q}}$. After that, we utilize a repositioning transformer[5] to predict the matching matrix and transformation $R,t$. This matching matrix and transformation are supervised by the Matching loss $L_{M}$ and Warping loss $L_{W}$ from [5]. To train the denoising module, we utilize the Gaussian transition kernel to add noise to the ground truth matching matrix $E^{0}$ to make $E^{t}$ at timestamp $t$. Then we exploit focal loss to train the denoising module $g_{\theta}$. We exploit the focal loss $L_{simple}$ (which is the denoising matching term in Eqn.15) to optimize $g_{\theta}$. The total loss is defined as:
@@
Figure 3: Overview of our framework training. Our framework includes a KPConv[9] backbone optimization and a denoising module optimization. The training detail of the denoising module is listed in Algorithm.2. We implement the denoising matching loss $L_{simple}$ by utilizing a focal loss.
@@ IV-B Rigid datasets: 3DMatch and 3DLoMatch
TABLE IV: This result of reverse sampling process in deterministic DDIM (i.e., $z_{t}=0$ in Eqn.14). $Backbone$ denotes the $E^{T}$ generated by the backbone’s output. $Gaussian$ denote the $E^{T}$ is sampled from the Gaussian white noise $N(0,1)^{N{\times}M}$.
@@ IV-E Integrating Correspondences to Non-rigid Registration
GeoTR 92.0 0.296 74.0 0.284
GeoTR. + PEAL 1-step 93.7 0.663 77.8 0.642
GeoTR. + DiffusionPCR 1-step 93.9 0.625 78.2 0.620
@@
GeoTR. + PEAL 5-step 94.0 2.131 78.5 2.074
GeoTR. + DiffusionPCR 5-step 94.4 1.939 80.0 1.964

### 17 Jan 2024 v4

@@ I Introduction
  • To the best of our knowledge, we are the first to utilize the diffusion model in the doubly stochastic matrix space for iteratively searching optimal matching matrix solutions using the reverse sampling process. This matching matrix diffusion model can be integrated as a module in any other 2D-2D, 2D-3D, and 3D-3D registration problems.
@@ III-B Preliminaries
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}=\prod_{s=1}^{t}\alpha_{s}$.
@@
Based on Eqn.3, Eqn.4, and Eqn.5 and utilize the Bayes rule, the denoising step can be derived as:
$\begin{split}&q(x_{t-1}|x_{t},x_{0})=\frac{q(x_{t}|x_{t-1},x_{0})q(x_{t-1}|x_{0})}{q(x_{t}|x_{0})}\\ &\propto N(x_{t-1};\underbrace{\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})x_{t}+\sqrt{\bar{\alpha}_{t-1}}(1-\alpha_{t})x_{0}}{1-\bar{\alpha}_{t}}}_{\mu_{q}(x_{t},x_{0})},\\ &\underbrace{\frac{(1-\alpha_{t})(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\textbf{I}}_{\Sigma_{q}(t)}).\end{split}$ (6)
Based on equation Eqn.2 and rewritten form of transition kernel $q(x_{t}|x_{t-1})=q(x_{t}|x_{t-1},x_{0})$ (due to Markov property), we can derive the ELBO:
$&logp(x_{0})\geq-L_{vb}=\\ &\underbrace{\mathbb{E}_{q(x_{1}|x_{0})}[logp_{\theta}(x_{0}|x_{1})]}_{reconstruction\ term}-\underbrace{D_{KL}(q(x_{T}|x_{0}||p(x_{T})))}_{prior\ matching\ term}\\ &-{\sum\limits_{T=2}^{T}}\underbrace{\mathbb{E}_{q(x_{t}|x_{0})}[D_{KL}(q(x_{t-1}|q_{x_{t}},x_{0}))||p_{\theta}(x_{t-1}|x_{t})]}_{denoising\ matching\ term}\end{aligned}$ (7)
By some derivations[43], the optimization of denoising matching term (in Eqn.7) boils down to learning a network $g_{\theta}(x_{t})$ to predict the ground truth $x_{0}$ (i.e., in $\mu_{q}(x_{t},x_{0})$) from an arbitrarily diffused version $x_{t}$:
@@
The images or point clouds obtained from sensors undergo scale transformation and perspective changes. The data of scenes exhibit symmetry, globally repeating, as well as phenomena like noise or non-rigid deformation. Matrix space can effectively represent these phenomena and has higher dimensions than low dimensions space,(e.g., SE(3) and $\mathbb{R}^{3N}/\mathbb{R}^{2N}$). Therefore, searching directly in the matching matrix space theoretically enables us to find better solutions. In the registration tasks, the many-to-one relationship between the matching matrix space and the warping operation space serves as motivation for us to propose searching in the matching matrix space:
Hypothesis 1.
If there is a network that can accurately predict the matching degree between $p_{i}$ and $q_{j}$ with its score $E_{ij}$, and if there is an efficient optimization algorithm capable of searching in the matrix space, then under these conditions, searching for the optimal solution in the match matrix space may yield superior results compared to direct search in low-dimensional spaces(e.g., SE(3) and $\mathbb{R}^{3N}/\mathbb{R}^{2N}$)
For instance, we can give an informal formulation of the rigid scene:
$\underset{E\in\mathcal{M}}{\min}\sum\limits_{i=1}^{N}\sum\limits_{j=1}^{M}E_{ij}*||W_{E}(p_{i})-q_{j}||_{2}$
$\leq\mathop{\min}\limits_{W\in SE(3)}\sum\limits_{i=1}^{N}\sum\limits_{j=1}^{M}\hat{E}_{ij}*||W(p_{i})-q_{j}||_{2}$ (10)
where $W_{E}$ is a warping that generated based on $E$ and $\hat{E}$ is a fixed matching matrix pre-computed.
Figure 2: Overview of our matching matrix diffusion model. $\bigoplus$ mean 3D point coordinates and position encoding are both utilized as input. $\bigotimes$ means only 3D point utilized. The input points $\hat{P},\hat{Q}$ along with their corresponding point features $F^{\hat{P}},F^{\hat{Q}}$ remain fixed throughout the entire reverse sampling process after being output from the KPConv backbone. These inputs undergo transformation through the warping operation and denoising transformer at each denoising step. The forward diffusion process is modeled by the Gaussian transition kernel $q(E^{t}|E^{t-1})$ which has a closed form $q(E^{t}|E^{0})$. The denoising model $g_{\theta}(E^{t})$ learns a reverse denoising gradient that points to the target solution $E^{0}$. During inference in the reverse sampling process, we utilize the predicted $\hat{E}_{0}$ and Eqn. (5,15) to sampling $E^{t-1}$.
@@ III-C Framework Overview
Our framework consists of a KPConv[9] feature backbone and one diffusion model[21]. The backbone takes source point clouds $P$ and target $Q$ as input and performs three downsamplings and two upsamplings to obtain the downsampled super points $\hat{P}$ and $\hat{Q}$, along with their associated features associate features $F^{\hat{P}}\in\mathbb{R}^{N{\times}d}$ and $F^{\hat{Q}}\in\mathbb{R}^{M{\times}d}$. The initilized matching matrix $E^{T}$ is sampled from white noise or computed by the inner product of $F^{\hat{P}}$ and $F^{\hat{Q}}$. We then employ the denoising module $g_{\theta}$ (III-E) to reverse sample the target matching matrix $E^{0}$.
@@ III-D Diffusion Model in Doubly Stochastic Matrix space
and it is defined only depend on $\alpha_{t}$,$\bar{\alpha}_{t}$,and $E^{0}$ (see Eqn.6).
@@
Optimization loss. Inspired by [46] and Eqn.8, we replace the ”denoising matching term” in ELBO (Eqn.7) by a simplified version:
$L_{simple}=-\mathbb{E}_{q(\hat{E}^{0})}\left[\sum\limits_{t-1}^{T}\left(\frac{T-t+1}{T}\right)\mathbb{E}_{q(E^{t}|\hat{E}^{0})}logp_{\theta}(\hat{E}^{0}|E^{t})\right]$ (16)
where the reweighting term $\frac{T-t+1}{T}$ means that the loss at time steps close to T are weighted less than earlier steps. We give a brief derivation about this replacement $\mathbb{E}_{q(x_{t}|x_{0})}[D_{KL}(q(x_{t-1}|q_{x_{t}},x_{0}))||p_{\theta}(x_{t-1}|x_{t})]\rightarrow\mathbb{E}_{q(E^{t}|\hat{E}^{0})}logp_{\theta}(\hat{E}^{0}|E^{t})$:
$&logp(E^{0})\geq\mathbb{E}_{E^{1:T}\sim q(E^{1:T}|E^{0})}\left[log\left(\frac{p_{\theta}(E^{0:T})}{q(E^{1:T}|E^{0})}\right)\right]\\ &=\mathbb{E}_{q}\left[log\left(\frac{p_{\theta}(E^{0:T})}{q(E^{1:T}|E^{0})}\right)\right]\\ &=\mathbb{E}_{q}\left[logp_{\theta}(E^{T})+log(\frac{p_{\theta}(E^{0}|E^{1})}{q(E^{1}|E^{0})})+\sum\limits_{t=2}^{T}log(\frac{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}p_{\theta}(E^{t-1}|E^{t})}}{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}q(E^{t}|E^{t-1})}})\right]\\ &=\mathbb{E}_{q}\left[logp_{\theta}(E^{T})+log(\frac{p_{\theta}(E^{0}|E^{1})}{q(E^{1}|E^{0})})\right]\\ &+\mathbb{E}_{q}\left[\sum\limits_{t=2}^{T}log\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\frac{q(E^{t-1}|E^{0})}{q(E^{t}|E^{0})}}\cdot\frac{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}q(E^{t-1}|E^{t},E^{0})p_{\theta}(E^{0}|E^{t})}}{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}q(E^{t-1}|E^{t},E^{0})}}\right)\right]\\ &=\mathbb{E}_{q}\left[S_{\theta}(E^{T},E^{1},E^{0})+\sum\limits_{t=2}^{T}log(p_{\theta}(E^{0}|E^{t}))\right]\end{aligned}$
where $S_{\theta}(E^{T},E^{1},E^{0})=log\left(\frac{p_{\theta}(E^{T})}{q(E^{T}|E^{0})}\right)+log(p_{\theta}(E^{0}|E^{1}))$ and the blue equality is from Eqn.13. Since we employ the denoising module $g_{\theta}(\cdot)$ to predict the target $E^{0}$, we can reparameterize the $p_{\theta}(E^{t-1}|E^{t})$ as:
$p_{\theta}(E^{t-1}|E^{t})\propto q(E^{t-1}|E^{t},E^{0})p_{\theta}(E^{0}|E^{t}).$ (17)
@@ III-E2 Weighted SVD: $\mathbf{soft\_procrustes(\cdot,\cdot,\cdot)}$
$H=\sum\limits_{(i,j)\in K}\tilde{E}(i,j)\hat{p}_{i}\hat{q}_{j}^{T},\ H=U{\Sigma}V^{T}$ (19)
@@
$\mathbf{t}=\frac{1}{|K|}\left({\sum\limits_{(i,\cdot)\in K}\ \hat{p}_{i}}-\mathbf{R}{\sum\limits_{(\cdot,j)\in K}}\hat{q}_{j}\right)$ (21)
@@ III-E3 Warping Function: $\mathbf{warping(\cdot,\cdot,\cdot)}$
In this paper, we utilize the rigid warping for both rigid and deformable registration cases to fastly demonstrate our idea since the deformation of adjacent frame point clouds in deformable scenes is not significant. With the predicted correspondences and their associated local rigid transformation assumptions, we can perform nearest neighbor interpolation in the predicted correspondences to interpolate the flow of any point in $\hat{P}$. For the deformation warping, we actually can build a deformation graph[48, 49] to conduct deformable warping[50, 13] for the source point cloud $\hat{P}$. We leave the deformable warping integration in future work.
@@ III-E4 Denoising Transformer: $f_{\theta}(\cdot,\cdot,\cdot,\cdot,\cdot,\cdot)$
Attention layer. Following [5], we integrate the rotary position encoding $\Theta(\cdot)$[5] in the attention layers. Specifically, in the self-attention, the $q,k,v$ are computed as:
@@
Since KPConv [9] is a translation invariant backbone, to avoid the asymmetric or globally repetitive ambiguities, we utilize the rotary positional encoding $\Theta(\cdot)$ (defined in [5]) as our position encoding in the Denoising Transformer and Matching function.
@@ III-F Framework Training
Our framework utilize a KPConv[9] branch to give the downsampled the superpoints $\hat{P}$ and $\hat{Q}$ and the asccociated features $F^{\hat{P}}$ and $F^{\hat{Q}}$. After that, we utilize a repositioning transformer[5] to predict the matching matrix and transformation $R,t$. This matching matrix and transformation are supervised by the Matching loss $L_{M}$ and Warping loss $L_{W}$ from [5]. To train the denoising module, we utilize the Gaussian transition kernel to add noise to the ground truth matching matrix $E^{0}$ to make $E^{t}$ at timestamp $t$. Then we exploit focal loss to train the denoising module $g_{\theta}$. We exploit the focal loss $L_{simple}$ (which is the denoising matching term in Eqn.16) to optimize $g_{\theta}$. The total loss is defined as:
@@
Figure 3: Overview of our framework training. Our framework includes a KPConv[9] featuren backbone optimization and a denoising module optimization. The training detail of the denoising module is listed in Algorithm.2. We implement the denoising matching loss $L_{simple}$ by utilizing a focal loss. After the model has been trained, only the KPConv[9] backbone and denoising module remain for use in the reverse sampling process, while the other components are discarded.
@@ IV-B Rigid datasets: 3DMatch and 3DLoMatch
TABLE IV: This result of reverse sampling process in deterministic DDIM (i.e., $z_{t}=0$ in Eqn.15). $Backbone$ denotes the $E^{T}$ generated by the backbone’s output. $Gaussian$ denote the $E^{T}$ is sampled from the Gaussian white noise $N(0,1)^{N{\times}M}$.
@@ IV-E Integrating Correspondences to Non-rigid Registration
GeoTR[4] 92.0 0.296 74.0 0.284
GeoTR. + PEAL[16] 1-step 93.7 0.663 77.8 0.642
GeoTR. + DiffusionPCR[34] 1-step 93.9 0.625 78.2 0.620
@@
GeoTR. + PEAL[16] 5-step 94.0 2.131 78.5 2.074
GeoTR. + DiffusionPCR[34] 5-step 94.4 1.939 80.0 1.964